Circular No. 5851

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J1008–57 Y. Tanaka, Institute of Space and Astronautical Science, on behalf of

the ASCA Team, communicates: “The hard x-ray transient pulsar GRO J1008–57 (IAUC 5836, 5838) was observed with ASCA between Aug. 5.57 and 6.17 UT. The source was detected with an average intensity of ∼ 20 mCrab in the range 1-10 keV. The source position was determined to be α = 10^{h}09^{m}46^{s}, δ = −58^{o}17′38′′ (equinox 2000.0; error radius 1′). The intensity was modulated with a period of 93.6 s, which is consistent with that reported by the Gamma Ray Observatory. The x-ray pulse profile has a double-peak structure with a well-defined, sharp intensity minimum and a less prominent secondary minimum. The spectrum is hard and can be represented by a power law with a photon index of ∼ 1 and a significant absorption measure of ∼ 1.3 × 10^{22} cm^{−2}. The energy flux is ∼ 6 × 10^{−10} erg cm^{−2} s^{−1} in the range 2-10 keV.”

GRO J0422+32 Tanaka also reports: “Following the notices of an optical flux increase

of GRO J0422+32 (IAUC 5842), we performed an x-ray observation of the source with ASCA on Aug. 26 UT. We found an x-ray source at the position of the previously identified optical counterpart of GRO J0422+32 with an intensity of ∼ 1 mCrab, or ∼ 2 × 10^{−11} erg cm^{−2} s^{−1} in the range 0.5-10 keV. The spectrum is apparently featureless and represented by a power law with a slope (photon index) of ∼ 1.6.”

D. W. Hoard and S. Wachter, University of Washington; and P. Garnavich, Dominion Astrophysical Observatory, communicate: “Further CCD photometry of the optical counterpart of the soft x-ray transient GRO J0422+32 has been obtained at the Manastash Ridge Observatory. The preliminary V magnitudes (± 0.06) are: Aug. 17.45 UT, 15.43; 18.45, 15.40; 19.44, 15.37; 22.43, 15.32; 23.43, 15.23; 24.38, 15.31.”

R CORONAE BOREALIS This variable has begun to fade in brightness, as indicated by the

following visual magnitude estimates: Aug. 5.89 UT, 6.0 (A. Pereira, Cabo da Roca, Portugal); 18.88, 6.0 (Pereira); 23.13, 6.1 (W. G. Dillon, Missouri City, TX); 27.15, 6.4 (Dillon); 27.87, 6.6 (Pereira); 28.83, 6.7 (H. Bengtsson, Hisings Backa, Sweden); 29.9, 6.9 (M. Tombelli, Montelupo, Italy).

1993 August 30 Daniel W. E. Green